OpenStudy (fluttershyk):

find the slope of the line through (-9, -10) and (-2, -5) -5/7 7/5 5/7 -7/5>>>my answer

OpenStudy (fluttershyk):

@rebeccaxhawaii @KamiBug

OpenStudy (er.mohd.amir):

slope=(y2-y1)/(x2-x1)

OpenStudy (fluttershyk):

-5/7

OpenStudy (anonymous):

\[\frac{-5--10}{-2--9}=\frac{5}{7} \]

OpenStudy (retireed):

It is always easier to visualize a line problem, if you try to graph it. First I used the table function to plot the points. The I calculated the line to fit the point. This is a very user friendly graphing calculating web site. If you are not allowed to use a calculator, just rough sketching a graph on paper can help visualize the question.

OpenStudy (fluttershyk):

ok

OpenStudy (retireed):



OpenStudy (fluttershyk):

ok